OpenStudy (anonymous):

How do you restrict the domain of h(x) = 3(x+2)^2-1 so that h^-1 would be a function?

OpenStudy (anonymous):

\[h(x)=3(x+2)^2-1\] has vertex at \[(-2,-1)\] and is a parabola that opens up, so it is decreasing on \[(-\infty,-2)\] and increasing on \[(-2,\infty)\]

OpenStudy (anonymous):

to make it one to one make sure it is always decreasing or increasing. so you can pick either interval above

OpenStudy (anonymous):

what does that symbol mean?

OpenStudy (anonymous):

inifinity \[\infty\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

and how would you write that as a restriction?

OpenStudy (anonymous):

\[x\geq -2\] would work